# Beyond Complete Shapes: A Quantitative Evaluation of 3D Shape Matching Algorithms

### v1

@@ 5.4.1 Challenge 1: Scaling of Shapes
Table 4: Scale Analysis: Correspondence prediction accuracy measured by Area-under-the-curve (AUC x100) for baselines on all three shape matching settings. Unsupervised ULRSSM in the full-to-full setting, supervised GeomFmaps in partial-to-full and DPFM in all settings exhibit a notable performance drop (underlined) in the default compared to the normalised scale setting. In the full-to-full setting, the methods automatically normalise the shapes, which is not possible in partial settings without additional information.
5.4.2 Challenge 2: Alignment of Shapes
Most existing partial correspondence datasets, like CP2P [3] and SHREC’16 [17], keep shapes in the same canonical pose during training and testing. This bias can prevent methods from generalising to shapes under random rigid transformations, see also [4]. As this assumption is often unrealistic in real-world settings, we introduce random rotations along one axis in our benchmark in an attempt towards a more realistic setting. In Fig. 6, we show the effects of rotation on the partial-to-full setting. While reliable intrinsic (and therefore rotation-invariant) features such as wave kernel signature (wks) exist, these are not robust to partiality. With that, XYZ coordinates are commonly used as input for partial settings, making the performance of learning-based methods (particularly unsupervised methods) depend strongly on initial alignment. We thus compare two different settings to probe existing methods for robustness to shape rotation:
@@ 5.4.2 Challenge 2: Alignment of Shapes
    Our default benchmark setting with rotation, where all shapes in the train and test sets have a random rotation around one axis (i.e. all pairs of shapes are in different poses). This is the most natural setting among the considered ones, but also most difficult.
@@
    without rotation, where the train and test are not rotated (i.e. all pairs of shapes are rigidly aligned, similar to SHREC’16 [17] in which most shapes are in a canonical pose).
A noticeable decline in performance is observed when rotation is introduced compared to when the shapes are aligned. Experiments with aligned shapes in training and testing outperform those with random rotation around one axis in all learning-based methods. This indicates that current methods benefit from biased datasets, which stems from the use of XYZ as input features. This presents a challenge for future research in the domain of partial shape matching involving rotated partial shapes. One important question is how to define rotation-invariant feature descriptors that are robust to partiality. We believe our benchmark may serve as an important contribution to tackling this challenge.
@@
Figure 6: Rotation evaluation for the train and test sets of BeCoS presented in Section 5.4.2 with (solid) and without (dashed) rotation. Unsupervised methods struggle the most with rotation, whereas supervised methods can better accommodate rotational changes in the data due to their supervision signals.
6 Discussion and Limitations
Our work highlights shortcomings of existing 3D shape matching datasets for partial shape matching. We propose novel ways for the procedural generation of more challenging and relevant partial shape matching instances that are particularly well-suited for the demands of modern machine learning. We hope that our work will foster the development of more robust, more accurate and more efficient algorithms for 3D shape matching. While we do not see direct negative societal impacts of our work, similar as in advances in other applied areas of machine learning, improved algorithms may enable more involved misuse scenarios (e.g. related to forgery, or surveillance).
The cross-category and cross-dataset correspondences of BeCoS are generated pairwise for the selected template shapes. Although they are of reasonable quality, they may exhibit small errors, especially for cross-category shapes that may give rise to correspondence ambiguity (e.g. hooves to paws). Errors may potentially accumulate if many pairwise ground truth correspondences are chained (i.e. for shapes that are distant in our spanning tree of shapes), which deteriorates the composed ground truth quality. This could potentially be improved by incorporating cycle-consistency or synchronisation techniques [31, 43, 9] in the process of generating ground truth correspondences.
@@ 7 Conclusion
While shape matching has made enormous progress, existing methods still struggle with the problem of partial matching. To address this shortcoming, we proposed a novel data generation framework to efficiently create multi-category partial-shape datasets in a procedural way. This framework allows the generation of a virtually infinite number of partial-to-partial shape matching instances, thereby accommodating the demands of modern machine learning methods. Using our proposed framework, we aggregate a set of commonly used non-rigid correspondence datasets to procure a benchmark for partial non-rigid shape correspondence. The quantitative performance of recent methods on this new benchmark suggests that the partial setting is very challenging and demands more attention from the research community.
@@
The supplementary material provides more information about our framework and the BeCoS benchmark. In Section 8, we show additional results on partial-to-partial (P2P) setting. In Section 9 we give an overview of the used evaluation metrics. In Section 10, we provide details on the procedure followed to create our framework. Additionally, we present details about the BeCoS benchmark in Section 11. We discuss the licenses of our framework and the assets used to create the BeCoS benchmark and our long-term preservation plan in Section 12. In Section 13, we provide information about the computation resources used for the baseline methods.
8 The BeCoS Benchmark: Additional Results on the Partial-to-Partial Setting
8.1 Alternative Energies
In the main paper, we use XYZ coordinates for the axiomatic methods SM-COMB and GC-PPSM for the partial-to-partial setting. While these methods work best with learned features (e.g. unsupervised features from ULRSSM), the latter features are not reliable for our challenging benchmark dataset (see Fig. 5 in the main paper for the performance of ULRSSM in the partial-to-full setting). In our case, we found that XYZ coordinates as feature descriptors are most suited for our benchmark. We compare the XYZ features to the original energy formulation used for SM-COMB [48] (a combination of bending and membrane energy) and the rotation-invariant SHOT features. We show the results in Fig. 7 for the geodesic distance comparison and Table 5 for the overlapping region prediction. Both show that XYZ as input features work best in the given setting. This is the case because in BeCoS the shapes are rotated only around the z-axis, which leaves the z-coordinate as reliable information. While this is not always a realistic setting, as it would not work for completely random rotations, we choose this setting to be able to evaluate the axiomatic methods in a (reasonably) meaningful manner. As these methods will benefit from better features, we hope to encourage further research in this direction.
@@ 8.1 Alternative Energies
XYZ 47.04 60.72
SM-COMB Original 23.83 24.52
Shot 24.52 34.35
XYZ 49.34 62.49
GC-PPSM Original 43.84 53.84
Shot 40.05 51.34
@@
8.2 Overlapping Region
In the default configuration of BeCoS in the partial-to-partial setting, shapes have an overlap between 10% and 90%. We show a histogram of the distribution of the overlapping region of each shape pair going from shape $\mathcal{X}$ to shape $\mathcal{Y}$ and vice-versa (see Fig. 8).
Figure 8: Partial-to-Partial histogram of overlapping region of BeCoS. The pairs that are below 10% and the ones above 90% are the ones that reached the maximum number of iterations $m=10$. Our framework produces partial-to-partial shapes, creating this histogram when the predefined parameters discussed in Sec 4 of the main paper are used.
9 Evaluation Metrics
For a better understanding of the used quantitative evaluation we give further insight about the following well-established quantitative measures.
Geodesic Error.
We evaluate the geodesic error of correspondences using the Princeton Protocol [33]. This evaluation entails normalising the geodesic error between the ground truth and the computed correspondences by the shape diameter, defined as the square root of the area of the target shape. For partial shapes, we use the square root of the area of their corresponding full shape. To handle the partial-to-partial matching paradigm, we set an infinite geodesic error for those unmatched vertices that should be matched and for those vertices that are matched but should not be, like in [24]. As we possibly have vertices with infinite geodesic error, the geodesic error curve in Fig. 5 in the main paper on the right does not necessarily reach 100%. By calculating the geodesic error only for the matched vertices, we get a metric that encompasses the overall matching performance.
Intersection over Union.
To evaluate the quality of the overlapping areas induced by a matching, we consider the Intersection over Union (IoU). To this end, we define $P\in\{0,1\}^{(|V|\times 1)}$ and $G\in\{0,1\}^{(|V|\times 1)}$ as vectors (predicted and ground truth, respectively), indicating which vertices are part of the matching. We quantitatively evaluate the overlapping region in the partial-to-partial setting using the intersection over union (IoU), calculated as $\text{IoU}=\frac{|P\cap G|}{|P\cup G|}$.
F1 Score.
To give further insight into the overlap prediction, we use the F1 score. It integrates precision ($\frac{\sum P\cap G}{\sum P}$) and recall ($\frac{\sum P\cap G}{\sum G}$) through the harmonic mean: F1 score = $\frac{2\cdot\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}$. Its possible value spans $[0,1]$ with $1$ indicating a perfect precision and recall.
@@ 10 Details on the Procedural Data Generation
10.1 Overlap Region for the Partial-to-Partial Setting
For the overlap computation, we rigidly align the shapes $\mathcal{X}$ and $\mathcal{Y}$ using the given correspondences to ensure that both partial shapes are semantically similar and perform the previously described partiality generation on both shapes. Subsequently, we quantify the overlap as the percentage of vertices from $\hat{\mathcal{X}}$ that are mapped to $\hat{\mathcal{Y}}$ and vice-versa. If both overlaps do not fall within a predefined range, the process iterates with new random camera poses until the overlap meets the specified criteria, or until a maximum of $m$ iterations is reached. In the latter case, the partial shapes whose overlap is the closest to the desired range are selected.
@@ 10.2 Correspondence Generation
This procedure is similar to the one used by [39], and we use FaceForm software with an academic license [28]:
@@
    We select a template for each dataset with cross-category correspondence and a template for each category in each dataset that doesn’t have cross-category correspondences. The template should ideally be in a neutral A-or T-pose for humanoid shapes and in a "upright" standing position for four-legged.
@@
    We build a tree connecting all templates in such a way that the semantic differences between two connected shape is small. This makes the morphing step easier. For example, we try to connect all cats from all datasets together (see the interactive graph in our project page).
@@
    For each edge in the tree, we load the two connected templates $A$ and $B$ in FaceForm using the "LoadGeom" module (see Fig. 9 (c)).
@@
    Some template shapes do not completely match the other template as some parts are missing (see Fig. 10). For these shapes we first cut the shapes, such that we get a matchable and an unmatchable part. For these template shapes, we only load the matchable part in FaceForm.
@@
    To get the correspondences between the two shapes, we find for every vertex $\hat{A}$ the closest triangle of template shape $B$. We then project every vertex of $\hat{A}$ on the closest triangle and return the barycentric coordinates. The correspondences of unmatchable parts are set to -1.
@@
    We repeat steps [1-7] for all edges in the tree, such that we have the connection between any two template shapes.
@@ 10.3 Correspondence Computation between a Shape Pair
Given any two shapes in the dataset, we compute the ground-truth correspondences between using the following steps: First we extract the path between those two shapes over the given template shapes. In the interactive graph in our project page, we show the chosen templates and their inter-connectivity. We then extract the correspondences between any two shapes on the graph either through the given datasets or by using our manually annotated correspondences. We then propagate the correspondences along the path.
@@
Figure 9: Usage of FaceForm to generate the cross-category and cross-dataset ground-truth correspondences. (a) Given two templates, the user needs to manually place corresponding landmarks on both templates. (b) Overlap of the template A (blue) and the morphed template B (green) on top of each other after running the "Wrapping" module. (c) The modules used to perform the morphing of template B to template A.
@@
Figure 10: Correspondence between pairs with un-matchable parts. The colour is transferred from the left shape to the right shape. The red parts are the parts on the right shape that cannot be semantically matched on the left. (a) and (b) The centaur shape is a special case that can be matched to both humanoid and four-legged animals. (c) The elephant’s tusks and ears cannot be matched to a sheep. (d), (e) and (f) The horns are not always matched between shapes. (g) The rhino’s horn is un-matchable to any other animal in BeCoS. (h) The cat’s long tail is not matchable to the buck’s small fluffy tail.

### v2

@@ 5.4 Ablation Studies
Table 4: Scale Analysis: Correspondence prediction accuracy measured by Area-under-the-curve (AUC x100) for baselines on all three shape matching settings. Unsupervised ULRSSM in the full-to-full setting, supervised GeomFmaps in partial-to-full and DPFM in all settings exhibit a notable performance drop (underlined) in the default compared to the normalised scale setting. The best performing method in each setting is in bold.
Scaling of Shapes. To illustrate the effects of scaling, we compare the performance of multiple shape matching methods using the default setting of BeCoS (as discussed in Sec. 3.2) and a normalised setting, where the surface area of each (full) shape in the benchmark is normalised to $1$. The results in Table 4 show that the performance drops when using the default setting for most methods. While full-to-full methods can easily handle the full-to-full variant of BeCoS under the default setting by normalising the surface area to $1$, resulting in similarly scaled shapes. This is non-trivial for the partial-to-full and partial-to-partial settings, as each partial shape represents an unknown percentage of its full counterpart, making the normalisation non-trivial. Using the normalised surface area under these two partial settings thus increases performance, especially for the supervised learning-based methods DPFM and GeomFmaps. On the other hand, the axiomatic and unsupervised methods generally perform poorly on BeCoS, making it hard to draw a conclusion about the methods’ performance in the normalised and default settings.
Alignment of Shapes. To demonstrate the impact of shape alignment on the performance of shape matching methods, we compare two different settings to probe existing methods for robustness to shape rotation on the partial-to-full setting:
@@
    With rotation: Our default realistic benchmark setting with rotation, where all shapes in the train and test sets have a random rotation around the $z$-axis (i.e. all pairs of shapes are in different poses).
@@
    Without rotation: The train and test are not rotated, and shapes are kept as present in their respective datasets (most shapes are in a canonical pose).
We select the partial-to-full methods for this ablation, as the partial-to-partial methods perform prohibitively poorly, resulting in a mostly random prediction accuracy. Many full-to-full methods use wave kernel signature features [aubry_wave_2011] as input, making them independent of rotations.
We show the results of our ablation in Fig. 6. A noticeable decline in performance is observed when rotation is introduced compared to when the shapes are kept as present in their datasets. This indicates that current methods benefit from biased datasets. The bias errantly favours XYZ features as input features when shapes are all aligned in a canonical pose. While reliable intrinsic (and therefore rotation-invariant) features such as wave kernel signature (wks) exist, these are not robust to partiality. This presents a challenge for future research in the domain of partial shape matching involving rotated partial shapes. One important question is how to define rotation-invariant feature descriptors that are robust to partiality. We believe our benchmark may serve as an important contribution to tackling this challenge.
@@
Figure 6: Rotation evaluation for the test set of BeCoS presented in Sec. 5.4 with (solid) and without (dashed) rotation. Unsupervised methods struggle the most with rotation, whereas supervised methods can better accommodate rotational changes in the data due to their supervision signals.
5.5 Custom Annotations: Left/Right Predictions
As a proof of concept, we show our framework’s ability to propagate custom annotations throughout the dataset. To this end, we provide a quantitative evaluation of the performance of shape matching methods in predicting left/right parts on our benchmark. In this experiment, we use the left/right accuracy as evaluation metric (see Sec. 5.1 for more details) and compare all baselines presented in Sec. 5.2. We present the results on our benchmark in Table 5. Similar to the dense correspondence prediction, the supervised methods perform best in this task. While unsupervised methods perform well in full-to-full shape matching, in the partial-to-full case, they perform poorly due to the XYZ coordinates as input (shapes are by default randomly rotated around the $x$-axis in our benchmark). Similarly, the partial-to-partial methods perform unconvincingly.
Methods Left/Right Accuracy ($\uparrow$)
F2F GeomFmaps (Sup) 88.09
DPFM (Sup) 89.05
DPFM (Unsup) 86.92
ULRSSM (Unsup) 82.13
Smooth Shells (Axio) 72.88
P2F GeomFmaps (Sup) 78.96
DPFM (Sup) 80.84
DPFM (Unsup) 51.24
ULRSSM (Unsup) 53.39
PFM (Axio) 52.85
P2P DPFM (Sup) 29.17
SM-COMB (Axio) 14.23
GC-PPSM (Axio) 16.82
Table 5: Chirality benchmark: While methods in the full-to-full setting and supervised methods in the partial-to-full setting perform well, unsupervised methods in the partial-to-full setting and all methods in the partial-to-partial setting perform poorly.
6 Discussion and Future Work
The cross-category and cross-dataset correspondences of BeCoS are generated pairwise for the selected template shapes. Although they are of good quality, for cross-category shapes they may give rise to correspondence ambiguity (e.g. hooves to paws). Ambiguities may potentially accumulate if many pairwise ground truth correspondences are chained (i.e. for shapes that are distant in our shape network of shapes), which deteriorates the composed ground truth quality (see left example in Figure 8 in the supplementary material). This could potentially be improved by incorporating cycle-consistency or synchronisation techniques [huang2013consistent, pachauri2013solving, bernard2021sparse] in the process of generating ground truth correspondences.
@@ 7 Conclusion
While several shape matching datasets exist, they are typically static and limited in size, which constrains their adaptability to diverse problem settings and custom use cases. To address this we connect multiple existing mesh-based datasets through manual cross-category annotations, enabling dense correspondence and annotation propagation across a diverse set of shapes. Building on this foundation, we introduced a scalable framework for generating unlimited training instances for partial shape correspondence, simulating realistic occlusions and semantic variability. As a concrete instantiation, we release a challenging non-rigid shape surface matching benchmark and provided a comprehensive evaluation of current state-of-the-art methods. Our findings highlight persistent challenges, particularly in partial shape matching, and underscore the need for more robust and generalizable solutions. We hope our benchmark and framework will foster future advancements in shape matching research.
Acknowledgments
We thank Paul Roetzer and Daniel Scholz for their insightful feedback. This work is supported by the ERC starting grant no. 101160648 (Harmony), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 458610525 and the ERC Advanced Grant grant no. 884679 (SIMULACRON). Open Access funding enabled and organized by Projekt DEAL.
[…]
Beyond Complete Shapes: A Benchmark for Quantitative Evaluation
of 3D Shape Surface Matching Algorithms
@@ Acknowledgments
The supplementary material provides more information about our framework and the BeCoS benchmark. In Sec. 8, we show additional results on the partial-to-partial (P2P) setting and qualitative results of custom annotation propagation. In Sec. 9, we provide details on the procedure followed to create our framework. Additionally, we present details about the BeCoS benchmark in Sec. 10. We discuss the licenses of our framework and the assets used to create the BeCoS benchmark and our long-term preservation plan in Sec. 11. In Sec. 12, we provide information about the computation resources used for the baseline methods.
8 The BeCoS Benchmark: Additional Results
8.1 Alternative Energies on the Partial-to-Partial Setting
In the main paper, we use XYZ coordinates for the axiomatic methods SM-COMB [roetzer2022scalable] and GC-PPSM [ehm2024partial] for the partial-to-partial setting. While these methods work best with learned features (e.g. unsupervised features from ULRSSM [cao2023unsupervised]), the latter features are not reliable for our challenging benchmark dataset (see Fig. 5 in the main paper for the performance of ULRSSM in the partial-to-full setting). In our case, we found that XYZ coordinates as feature descriptors are most suited for our benchmark. We compare the XYZ features to the original energy formulation used for SM-COMB (a combination of bending and membrane energy) and the rotation-invariant SHOT features. We show the results in Fig. 7 for the geodesic distance comparison and Table 6 for the overlapping region prediction. Both show that XYZ as input features work best in the given setting. This is the case because in BeCoS the shapes are rotated only around the z-axis, which leaves the z-coordinates as reliable information. While this is not always a realistic setting, as it would not work for completely random rotations, we choose this setting to be able to evaluate the axiomatic methods in a (reasonably) meaningful manner. As these methods will benefit from better features, we hope to encourage further research in this direction.
@@ 8.1 Alternative Energies on the Partial-to-Partial Setting
XYZ 46.91 60.61
SM-COMB Original 23.81 33.78
Shot 24.47 34.29
XYZ 49.24 62.42
GC-PPSM Original 43.69 53.72
Shot 39.89 51.20
@@
8.2 Custom Annotation Propagation
We show more qualitative results of Left/Right annotation propagation using our framework in Fig. 8. We also include failure cases of the annotation propagation. This propagation relies on the manually created dense correspondence between templates in our framework; small errors caused by the manual annotation get amplified throughout the propagation and cause unwanted artefacts, especially in regions like the tails of animals.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Our framework allows for the propagation of custom annotations throughout all shapes. We show the case of left/right annotations. The Left/Right annotations exhibit errors in thin parts of the shapes (e.g. tails of animals).
@@ 9.1 Correspondence Generation
This procedure is similar to the one used by [magnet2022smooth], and we use the FaceForm software with an academic license [faceform]. The steps of our procedure can be summarised as follows:
@@
    We select a template for each dataset with cross-category correspondence and a template for each category in each dataset that does not have cross-category correspondences. The template should ideally be in a neutral A-or T-pose for humanoid shapes and in an "upright" standing position for four-legged animals.
@@
    We build a shape network connecting all templates in such a way that the semantic differences between two connected shapes are small. This makes the morphing step easier. For example, we try to connect all cats from all datasets together (see the interactive graph in our project page).
@@
    For each edge in the shape network, we load the two connected templates $A$ and $B$ in FaceForm using the "LoadGeom" module (see Fig. 10 (c)).
@@
    Some template shapes do not completely match the other template as some parts are missing (see Fig. 11). For these shapes, we first cut the shapes so that we get a matchable and a semantic missing part. For these template shapes, we only load the matchable part in FaceForm.
@@
    To get the correspondences between the two shapes, we find for every vertex $\hat{A}$ the closest triangle of template shape $B$. We then project every vertex of $\hat{A}$ on the closest triangle and return the barycentric coordinates. The correspondences of semantic missing parts are set to $-1$.
@@
    Once steps [3-7] are done for all edges in the tree, we are able to generate the correspondences between any two input shapes.
@@ 9.2 Correspondence Computation between a Shape Pair
Given any two shapes in the dataset, we compute the ground-truth correspondences between them using the following steps: First, we extract the path between those two shapes over the given template shapes. In the interactive shape network on our project page, we show the chosen templates and their inter-connectivity. We then extract the correspondences between any two shapes in the shape network either through the given datasets or by using our manually annotated correspondences. We then propagate the correspondences along the path.
9.3 Overlap Region for the P2P Setting
For the overlap computation, we rigidly align the shapes $\mathcal{X}$ and $\mathcal{Y}$ using the given correspondences to ensure that both partial shapes are semantically similar and perform the in the main paper described partiality generation on both shapes. Subsequently, we quantify the overlap as the percentage of vertices from $\hat{\mathcal{X}}$ that are mapped to $\hat{\mathcal{Y}}$ and vice-versa ($\hat{\mathcal{X}}$ and $\hat{\mathcal{Y}}$ are the partial shape generated using ray casting from shapes $\mathcal{X}$ and $\mathcal{Y}$, respectively). If both overlaps do not fall within a predefined range, the process iterates with new random camera poses until the overlap meets the specified criteria, or until a maximum of $m$ iterations is reached. In the latter case, the partial shapes whose overlap is the closest to the desired range are selected.
In the BeCoS default configuration for the partial-to-partial setting, shapes have an overlap between 10% and 90%. We show a histogram of the distribution of the overlapping region of each shape pair going from shape $\mathcal{X}$ to shape $\mathcal{Y}$ and vice-versa (see Fig. 9).
Figure 9: Partial-to-Partial histogram of overlapping region of BeCoS. The pairs that are below 10% and the ones above 90% are the ones that reached the maximum number of iterations $m=10$. Our framework produces partial-to-partial shapes, creating this histogram when the predefined parameters discussed in Sec. 4 of the main paper are used.
@@ 9.3 Overlap Region for the P2P Setting
Refer to caption Refer to caption
(c) (d)
Figure 10: Usage of FaceForm to generate the cross-category and cross-dataset ground-truth correspondences. (a) Given two templates, we manually place corresponding landmarks on both templates. (b) Overlap of the template A (blue) and the morphed template B (green) on top of each other after running the "Wrapping" module. (c) and (d) We show a non-isometric deformation between a puma and rhino shape. Note that the rhino is missing its horn which is considered a missing semantic part and we only match the rhino shape without its horn to the puma. (e) The modules used to perform the morphing of template B to template A.
@@
Figure 11: Correspondence between pairs with semantic missing parts. The colour is transferred from the left shape to the right shape. The red parts are the parts that cannot be semantically matched between the two shapes. (a) and (b) The centaur shape is a special case that can be matched to both humanoid and four-legged animals. (c) The elephant’s tusks and ears cannot be matched to a sheep. (d), (e) and (f) The horns are not always matched between shapes. (g) The rhino’s horn is not semantically matched to any other animal in BeCoS. (h) The cat’s long tail is not matched to the buck’s small fluffy tail.